OpenStudy (anonymous):

Let u = <-8, -9>. Find 7u.

jimthompson5910 (jim_thompson5910):

multiply every component by 7

jimthompson5910 (jim_thompson5910):

u = <-8, -9> 7*u = 7*<-8, -9> 7*u = <7*(-8), 7*(-9)> 7u = <-56, -63>